# Deep Learning for Camera Calibration and Beyond: A Survey

### v2

@@ Abstract
Camera calibration involves estimating camera parameters to infer geometric features from captured sequences, which is crucial for computer vision and robotics. However, conventional calibration is laborious and requires dedicated collection. Recent efforts show that learning-based solutions have the potential to be used in place of the repeatability works of manual calibrations. Among these solutions, various learning strategies, networks, geometric priors, and datasets have been investigated. In this paper, we provide a comprehensive survey of learning-based camera calibration techniques, by analyzing their strengths and limitations. Our main calibration categories include the standard pinhole camera model, distortion camera model, cross-view model, and cross-sensor model, following the research trend and extended applications. As there is no benchmark in this community, we collect a holistic calibration dataset that can serve as a public platform to evaluate the generalization of existing methods. It comprises both synthetic and real-world data, with images and videos captured by different cameras in diverse scenes. Toward the end of this paper, we discuss the challenges and provide further research directions. To our knowledge, this is the first survey for the learning-based camera calibration (spanned 8 years). The summarized methods, datasets, and benchmarks are available and will be regularly updated at https://github.com/KangLiao929/Awesome-Deep-Camera-Calibration.
@@
Camera calibration is a fundamental and indispensable field in computer vision and it has a long research history [1, 2, 3, 4], tracing back to around 60 years ago[5]. The first step for many vision and robotics tasks is to calibrate the intrinsic (image sensor and distortion parameters) and/or extrinsic (rotation and translation) camera parameters, ranging from computational photography, and multi-view geometry, to 3D reconstruction. In terms of the task type, there are different techniques to calibrate the standard pinhole camera, fisheye lens camera, stereo camera, light field camera, event camera, and LiDAR-camera system, etc. Figure 1 shows the popular calibration objectives, models, and extended applications in camera calibration.
Traditional methods for camera calibration generally depend on hand-crafted features and model assumptions. These methods can be broadly divided into three categories. The most prevalent one involves using a known calibration target (e.g., a checkerboard) as it is deliberately moved in the 3D scene [6, 7, 8]. Then, the camera captures the target from different viewpoints and the checkerboard corners are detected for calculating the camera parameters. However, such a procedure requires cumbersome manual interactions and it cannot achieve automatic calibration “in the wild”. To pursue better flexibility, the second category of camera calibration, i.e., the geometric-prior-based calibration has been largely studied [9, 10, 11, 12]. To be specific, the geometric structures are leveraged to model the 3D-2D correspondence in the scene, such as lines and vanishing points. However, this type of method heavily relies on structured man-made scenes containing rich geometric priors, leading to poor performance when applied to general environments. The third category is self-calibration[13, 14, 15]. Such a solution takes a sequence of images as inputs and estimates the camera parameters using multi-view geometry. The accuracy of self-calibration, however, is constrained by the limits of the feature detectors, which can be influenced by diverse lighting conditions and textures. The above parametric models are based on the physical interpretation of camera geometry. While they are user-friendly, they tend to be tailored for specific camera models and might not offer optimal accuracy. Instead, non-parametric models [16, 17, 18] link each image pixel to its 3D observation ray, eliminating the constraints of parametric models.
@@
Recently, deep learning has brought new inspirations to camera calibration and its applications. Learning-based methods achieve state-of-the-art performances on various tasks with higher efficiency. In particular, diverse deep neural networks (DNNs) have been developed, such as convolutional neural networks (CNNs), generative adversarial networks (GANs), PointNet, and vision transformers (ViTs), of which the high-level semantic features show more powerful representation capability compared with the hand-crafted features. Moreover, diverse learning strategies have been exploited to boost the geometric perception of neural networks. Learning-based methods offer a flexible and end-to-end camera calibration solution, without manual interventions or calibration targets, which sets them apart from traditional methods. Furthermore, some of these methods achieve camera model-free and label-free calibration, showing promising and meaningful applications.
@@
In this paper, we provide a comprehensive and in-depth overview of recent advances in learning-based camera calibration, covering over 100 papers. We also discuss potential directions for further improvements and examine various types of cameras and targets. To facilitate future research on different topics, we categorize the current solutions according to calibration objectives and applications. In addition to fundamental parameters such as focal length, rotation, and translation, we also provide detailed reviews for correcting image distortion (radial distortion and rolling shutter distortion), estimating cross-view mapping, calibrating camera-LiDAR systems, and other applications. Such a trend follows the development of cameras and market demands for virtual reality, autonomous driving, neural rendering, etc.
To our best knowledge, this is the first survey of the learning-based camera calibration and its extended applications, it has the following unique contributions. (1) Our work mainly follows recent advances in deep learning-based camera calibration. In-depth analysis and discussion in various aspects are offered, including publications, network architecture, loss functions, datasets, evaluation metrics, learning strategies, implementation platforms, etc. The detailed information of each literature is listed in Table I. (2) Despite the calibration algorithm, we comprehensively review the classical camera models and their extended models. In particular, we summarize the redesigned calibration objectives in deep learning since some traditional calibration objectives are verified to be hard to learn by neural networks. (3) We collect a dataset containing images and videos captured by different cameras in different environments, which can serve as a platform to evaluate the generalization of existing methods. (4) We discuss the open challenges of learning-based camera calibration and propose some future directions to provide guidance for further research in this field. (5) An open-source repository is created that provides a taxonomy of all reviewed works and benchmarks. The repository will be updated regularly in https://github.com/KangLiao929/Awesome-Deep-Camera-Calibration.
@@ 2.1 Learning Paradigm
Regression-based Calibration Given an uncalibrated input, the regression-based calibration first extracts the high-level semantic features using stacked convolutional layers. Then, the fully connected layers aggregate the semantic features and form a vector of the estimated calibration objective. The regressed parameters are used to conduct subsequent tasks such as distortion rectification, image warping, camera localization, etc. This paradigm is the earliest and has a dominant role in learning-based camera calibration and its applications. All the first works in various objectives, e.g., intrinsics: Deepfocal [24], extrinsic: PoseNet [25], radial distortion: Rong et al. [26], rolling shutter distortion: URS-CNN [27], homography matrix: DHN [28], hybrid parameters: Hold-Geoffroy et al. [29], camera-LiDAR parameters: RegNet [30] have been achieved with this paradigm.
Reconstruction-based Calibration On the other hand, the reconstruction-based calibration paradigm discards the parameter regression and directly learns the pixel-level mapping function between the uncalibrated input and target, inspired by the conditional image-to-image translation [31] and dense visual perception[32, 33]. The reconstructed results are then calculated for the pixel-wise loss with the ground truth. In this regard, most reconstruction-based calibration methods [34, 35, 36, 37] design their network architecture based on the fully convolutional network such as U-Net[38]. Specifically, an encoder-decoder network, with skip connections between the encoder and decoder features at the same spatial resolution, progressively extracts the features from low-level to high-level and effectively integrates multi-scale features. At the last convolutional layer, the learned features are aggregated into the target channel, reconstructing the calibrated result at the pixel level.
In contrast to the regression-based paradigm, the reconstruction-based paradigm does not require the label of diverse camera parameters. Besides, the imbalance loss problem can be eliminated since it only optimizes the photometric loss of calibrated results. Therefore, the reconstruction-based paradigm enables a blind camera calibration without a strong camera model assumption.
@@ 2.2 Learning Strategies
Supervised Learning Most learning-based camera calibration methods train their networks with the supervised learning strategy, from the classical methods [24, 25, 28, 39, 26, 40] to the state-of-the-art methods [41, 42, 43, 44, 45]. In terms of the learning paradigm, this strategy supervises the network with the ground truth of the camera parameters (regression-based paradigm) or paired data (reconstruction-based paradigm). In general, they synthesize the training dataset from other large-scale datasets, under the random parameter/transformation sampling and camera model simulation. Some recent works [46, 47, 48, 49] establish their training dataset using a real-world setup and label the captured images with manual annotations, thereby fostering advancements in this research domain.
Semi-Supervised Learning Training the network using an annotated dataset under diverse scenarios is an effective learning strategy. However, human annotation can be prone to errors, leading to inconsistent annotation quality or the inclusion of contaminated data. Consequently, increasing the training dataset to improve performance can be challenging due to the complexity and cost of constructing the dataset. To address this challenge, SS-WPC[50] proposes a semi-supervised method for correcting portraits captured by a wide-angle camera. It employs a surrogate task (segmentation) and a semi-supervised method that utilizes direction and range consistency and regression consistency to leverage both labeled and unlabeled data.
@@
Unsupervised Learning Unsupervised learning, commonly referred to as unsupervised machine learning, analyzes and groups unlabeled datasets using machine learning algorithms. UDHN [52] is the first work for a cross-view camera model using unsupervised learning, which estimates the homography matrix of a paired image without the projection labels. By reducing a pixel-wise intensity error that does not require ground truth data, UDHN [52] outperforms previous supervised learning techniques. While preserving superior accuracy and robustness to fluctuation in light, the proposed unsupervised algorithm can also achieve faster inference time. Inspired by this work, increasing more methods leverage the unsupervised learning strategy to estimate the homography such as CA-UDHN [53], BaseHomo [54], HomoGAN[55], and Liu et al. [56]. Besides, UnFishCor [57] frees the demands for distortion parameters and designs an unsupervised framework for the wide-angle camera.
Self-supervised Learning Robotics is where the phrase “self-supervised learning” first appears, as training data is automatically categorized by utilizing relationships between various input sensor signals. Compared to supervised learning, self-supervised learning leverages input data itself as the supervision. Many self-supervised techniques are presented to learn visual characteristics from massive amounts of unlabeled photos or videos without the need for time-consuming and expensive human annotations. SSR-Net [58] presents a self-supervised deep homography estimation network, which relaxes the need for ground truth annotations and leverages the invertibility constraints of homography. To be specific, SSR-Net [58] utilizes the homography matrix representation in place of other approaches’ typically-used 4-point parameterization, to apply the invertibility constraints. SIR [59] devises a brand-new self-supervised camera calibration pipeline for wide-angle image rectification, based on the principle that the corrected results of distorted images of the same scene taken with various lenses need to be the same. With self-supervised depth and pose learning as a proxy aim, Fang et al. [60] present to self-calibrate a range of generic camera models from raw video, offering for the first time a calibration evaluation of camera model parameters learned solely via self-supervision.
Reinforcement Learning Instead of aiming to minimize at each stage, reinforcement learning can maximize the cumulative benefits of a learning process as a whole. To date, DQN-RecNet [61] is the first and only work in camera calibration using reinforcement learning. It applies a deep reinforcement learning technique to tackle the fisheye image rectification by a single Markov Decision Process, which is a multi-step gradual calibration scheme. In this situation, the current fisheye image represents the state of the environment. The agent, Deep Q-Network [62], generates an action that should be executed to correct the distorted image.
@@
DHN[28] | RSSW | Cross-View | VGG | loss | MS-COCO[70] | MSE | SL | Caffe | ✓ | |
URS-CNN [27] | CVPR | Distortion | CNNs | loss | Sun[73], Oxford[74], Zubud[75], LFW[76] | PSNR, RMSE | SL | Torch | ✓ | |
@@
DeepFM[82] | ECCV | Cross-View | ResNet | loss | T&T[83], KITTI[77], 1DSfM[63] | F-score, Mean | SL | PyTorch | ✓ | |
2019 | Lopez et al. [88] | CVPR | Distortion | DenseNet | Bearing loss | SUN360[78] | MSE | SL | PyTorch | |
UprightNet [89] | ICCV | Standard | U-Net | Geometry loss | InteriorNet[90], ScanNet[91], SUN360[78] | Mean error | SL | PyTorch | ||
@@
SSR-Net [58] | PRL | Cross-View | ResNet | loss | MS-COCO[70] | MAE | SSL | PyTorch | ✓ | |
@@
Lee et al. [109] | ECCV | Standard | PointNet + CNNs | Cross-entropy loss | Google Street View[110], HLW[66] | Mean error, AUC[95] | SL | - | ||
@@
SRHEN [136] | ACM-MM | Cross-View | CNNs | loss | MS-COCO [70], SUN397 [78] | MACE | SL | - | ✓ | |
@@
DLKFM [146] | CVPR | Cross-View | Siamese-Net | loss | MS-COCO[70], Google Earth, Google Map | MSE | SL | TensorFlow | ✓ | |
@@
IHN [165] | CVPR | Cross-View | Siamese-Net | loss | MS-COCO[70], Google Earth, Google Map | MACE | SL | PyTorch | ✓ | |
@@
DiffPoseNet [170] | CVPR | Standard | CNNs + LSTM | loss | TartanAir[171], KITTI[77], TUM-RGBD[172] | PEE, AEE[173] | SSL | PyTorch | ||
DXQ-Net [179] | arXiv | Cross-Sensor | CNNs + RNNs | , geodesic loss | KITTI[77], KITTI-360[157] | MSE | SL | PyTorch | ✓ | |

### v3

@@ Abstract
Camera calibration involves estimating camera parameters to infer geometric features from captured sequences, which is crucial for computer vision and robotics. However, conventional calibration is laborious and requires dedicated collection. Recent efforts show that learning-based solutions have the potential to be used in place of the repeatability works of manual calibrations. Among these solutions, various learning strategies, networks, geometric priors, and datasets have been investigated. In this paper, we provide a comprehensive survey of learning-based camera calibration techniques, by analyzing their strengths and limitations. Our main calibration categories include the standard pinhole camera model, distortion camera model, cross-view model, and cross-sensor model, following the research trend and extended applications. As there is no unified benchmark in this community, we collect a holistic calibration dataset that can serve as a public platform to evaluate the generalization of existing methods. It comprises both synthetic and real-world data, with images and videos captured by different cameras in diverse scenes. Toward the end of this paper, we discuss the challenges and provide further research directions. To our knowledge, this is the first survey for the learning-based camera calibration (spanned 10 years). The summarized methods, datasets, and benchmarks are available and will be regularly updated at https://github.com/KangLiao929/Awesome-Deep-Camera-Calibration.
@@
Camera calibration is a fundamental and indispensable field in computer vision and it has a long research history [1, 2, 3, 4, 5], tracing back to around 60 years ago[6]. The first step for many vision and robotics tasks is to calibrate the intrinsic (image sensor and distortion parameters) and/or extrinsic (rotation and translation) camera parameters, ranging from computational photography, and multi-view geometry, to 3D reconstruction. In terms of the task type, there are different techniques to calibrate the standard pinhole camera, fisheye lens camera, stereo camera, light field camera, event camera, and LiDAR-camera system, etc. Figure 1 shows the popular calibration objectives, models, and extended applications in camera calibration.
Traditional methods for camera calibration generally depend on hand-crafted features and model assumptions. These methods can be broadly divided into three categories. The most prevalent one involves using a known calibration target (e.g., a checkerboard) as it is deliberately moved in the 3D scene [7, 8, 9]. Then, the camera captures the target from different viewpoints and the checkerboard corners are detected for calculating the camera parameters. However, such a procedure requires cumbersome manual interactions and it cannot achieve automatic calibration “in the wild”. To pursue better flexibility, the second category of camera calibration, i.e., the geometric-prior-based calibration has been largely studied [10, 11, 12, 13]. To be specific, the geometric structures are leveraged to model the 3D-2D correspondence in the scene, such as lines and vanishing points. However, this type of method heavily relies on structured man-made scenes containing rich geometric priors, leading to poor performance when applied to general environments. The third category is self-calibration[14, 15, 16]. Such a solution takes a sequence of images as inputs and estimates the camera parameters using multi-view geometry. The accuracy of self-calibration, however, is constrained by the limits of the feature detectors, which can be influenced by diverse lighting conditions and textures. The above parametric models are based on the physical interpretation of camera geometry. While they are user-friendly, they tend to be tailored for specific camera models and may not offer optimal accuracy. Instead, non-parametric models [17, 18, 19] link each image pixel to its 3D observation ray, eliminating the constraints of parametric models.
@@
Recently, deep learning has brought new inspirations to camera calibration and its applications. Learning-based methods achieve state-of-the-art performances on various tasks with higher efficiency. Particularly, diverse deep neural networks have been developed, such as convolutional neural networks (CNNs), generative adversarial networks (GANs), vision transformers (ViTs), and diffusion models, of which the high-level semantic features show more powerful representation capability compared with the hand-crafted features. Moreover, diverse learning strategies have been exploited to boost the geometric perception of networks. Learning-based methods offer a flexible and end-to-end camera calibration solution, without manual interventions or calibration targets, which sets them apart from traditional methods. Furthermore, some of these methods achieve camera model-free and label-free calibration, showing promising and meaningful applications.
@@
In this paper, we provide a comprehensive and in-depth overview of recent advances in learning-based camera calibration, covering over 200 related papers. We also discuss potential directions for further improvements and examine various types of cameras and targets. To facilitate future research on different topics, we categorize the current solutions according to calibration objectives and applications. In addition to fundamental parameters such as focal length, rotation, and translation, we also provide detailed reviews for correcting image distortion (radial distortion and rolling shutter distortion), estimating cross-view mapping, calibrating camera-LiDAR systems, and other applications. Such a trend follows the development of cameras and market demands for virtual reality, autonomous driving, neural rendering, etc.
To our best knowledge, this is the first survey of the learning-based camera calibration and its extended applications, it has the following unique contributions. (1) Our work mainly follows recent advances in deep learning-based camera calibration. In-depth analysis and discussion in various aspects are offered, including publications, network architecture, loss functions, datasets, evaluation metrics, learning strategies, implementation platforms, etc. The detailed information of each literature is listed in Table I and Table II. (2) Despite the calibration algorithm, we comprehensively review the classical camera models and their extended models. In particular, we summarize the redesigned calibration objectives in deep learning since some traditional calibration objectives are verified to be hard to learn by neural networks. (3) We collect a dataset containing images and videos captured by different cameras in different environments, which can serve as a platform to evaluate the generalization of existing methods. (4) We discuss the open challenges in learning-based camera calibration and propose future directions to guide further research in this field. (5) An open-source repository is created that provides a taxonomy of all reviewed works and benchmarks. The repository will be updated regularly in https://github.com/KangLiao929/Awesome-Deep-Camera-Calibration.
@@ 2.1 Learning Paradigm
Regression-based Calibration Given an uncalibrated input, the regression-based calibration extracts the high-level semantic features using stacked convolutional layers. Then, the fully connected layers aggregate the semantic features and form a vector of the estimated calibration objective. The regressed parameters are used to conduct subsequent tasks such as distortion rectification, image warping, camera localization, etc. This paradigm is the earliest and has a dominant role in learning-based camera calibration. All the first works in various objectives, e.g., intrinsics: Deepfocal [25], extrinsic: PoseNet [26], radial distortion: Rong et al. [27], rolling shutter distortion: URS-CNN [28], homography: DHN [29], hybrid parameters: Hold-Geoffroy et al. [30], camera-LiDAR parameters: RegNet [31] have been achieved with this paradigm.
Reconstruction-based Calibration On the other hand, the reconstruction-based calibration paradigm discards the parameter regression and directly learns the pixel-level mapping function between the uncalibrated input and target, inspired by the conditional image-to-image translation [32] and dense visual perception[33, 34]. The reconstructed results are then calculated for the pixel-wise loss with the ground truth. In this regard, most reconstruction-based calibration methods [35, 36, 37, 38] design their network architecture based on the fully convolutional network such as U-Net[39]. Specifically, an encoder-decoder network, with skip connections between the encoder and decoder features at the same spatial resolution, progressively extracts the features from low-level to high-level and effectively integrates multi-scale features. At the last convolutional layer, the learned features are aggregated into the target channel, forming the calibrated result or calibration representation at the pixel level. Recent works also explore harnessing the powerful generation ability of the diffusion model to help reconstruct the calibration targets [40, 41, 42, 43].
In contrast to the regression-based paradigm, the reconstruction-based paradigm does not require the label of diverse camera parameters during training. Besides, the imbalance loss problem can be eliminated since it only optimizes the photometric loss of calibrated results or calibration representations. Therefore, the reconstruction-based paradigm enables a blind camera calibration without a strong camera model assumption [17, 18, 19].
@@ 2.2 Learning Strategies
Supervised Learning Most learning-based camera calibration methods train their networks with the supervised learning strategy, from the classical methods [25, 26, 29, 44, 27, 45] to the state-of-the-art methods [46, 47, 48, 49, 50, 51]. In terms of the learning paradigm, this strategy supervises the network with the ground truth of the camera parameters (regression-based paradigm) or paired data (reconstruction-based paradigm). In general, they synthesize the training dataset from other large-scale datasets, under the random parameter/transformation sampling and camera model simulation. Some recent works [52, 53, 54, 55] establish their training dataset using a real-world setup and label the captured images with manual annotations, thereby fostering advancements in this research domain.
Semi-Supervised Learning Training the network using an annotated dataset under diverse scenarios is an effective learning strategy. However, human annotation can be prone to errors, leading to inconsistent annotation quality or the inclusion of contaminated data. Consequently, increasing the dataset to improve performance can be challenging due to the complexity and construction cost. To address this challenge, SS-WPC[56] proposes a semi-supervised method for correcting portraits captured by a wide-angle camera. It employs a surrogate segmentation task and a semi-supervised method that utilizes direction and range consistency and regression consistency to leverage both labeled and unlabeled data.
@@
Unsupervised Learning Unsupervised learning analyzes and groups unlabeled datasets using machine learning algorithms. UDHN [58] is the first work for the cross-view camera model using unsupervised learning, which estimates the homography matrix of a paired image without labels. By reducing a pixel-wise error that does not require ground truth data, UDHN [58] outperforms previous supervised learning techniques. While preserving superior accuracy and robustness to light fluctuations, it can also achieve faster inference time. Inspired by this work, increasing more methods leverage the unsupervised learning strategy to estimate the cross-view mapping such as CA-UDHN [59], BaseHomo [60], HomoGAN[61], and Liu et al. [62]. Besides, UnFishCor [63] frees the demands for distortion parameters and designs an unsupervised framework for the wide-angle camera.
Self-supervised Learning Robotics is where the phrase “self-supervised learning” first appears, as training data is automatically categorized by utilizing relationships between various input sensor signals [64]. Compared to supervised learning, self-supervised learning leverages input data itself as the supervision. Many self-supervised techniques are presented to learn visual characteristics from massive amounts of unlabeled photos or videos without the need for time-consuming and expensive human annotations. SSR-Net [65] presents a self-supervised deep homography estimation network, which relaxes the need for ground truth annotations and leverages the invertibility constraints of homography. To be specific, SSR-Net [65] utilizes the homography matrix representation in place of other approaches’ typically-used 4-point parameterization, to apply the invertibility constraints. SIR [66] devises a brand-new self-supervised pipeline for wide-angle image rectification, based on the principle that the corrected results of distorted images of the same scene captured by various lenses, should obtain identical rectified results. Using self-supervised depth and pose learning as a proxy task, Fang et al. [67] propose a method for self-calibrating a range of generic camera models from raw video, providing the first calibration evaluation of camera model parameters learned entirely through self-supervision.
Reinforcement Learning Instead of aiming to minimize at each stage, reinforcement learning can maximize the cumulative benefits of a learning process as a whole. To date, DQN-RecNet [68] is the first work in camera calibration using reinforcement learning. It applies a deep reinforcement learning technique to tackle the fisheye image rectification by a single Markov Decision Process, which is a multi-step gradual calibration scheme. In this situation, the current fisheye image represents the state of the environment. The agent, Deep Q-Network [69], generates an action that should be executed to correct the distorted image.
@@
2017 | URS-CNN [28] | CVPR | Distortion | CNNs | loss | Sun[77], Oxford[78], Zubud[79], LFW[80] | PSNR, RMSE | SL | Torch | ✓ |
@@
2019 | UprightNet [86] | ICCV | Standard | U-Net | Geometry loss | InteriorNet[87], ScanNet[88], SUN360[81] | Mean error | SL | PyTorch | |
Lopez et al. [93] | CVPR | Distortion | DenseNet | Bearing loss | SUN360[81] | MSE | SL | PyTorch | ||
@@
2020 | Lee et al. [104] | ECCV | Standard | PointNet+CNNs | Cross-entropy loss | Google Street View[105], HLW[73] | Mean error, AUC[106] | SL | - | |
@@
DiffPoseNet [149] | CVPR | Standard | CNNs+LSTMs | loss | TartanAir[150], KITTI[95], TUM RGB-D[151] | PEE, AEE[152] | SSL | PyTorch | ||
CCS-Net [159] | RAL | Distortion | U-Net | loss | TUM RGB-D[151] | MAE, RPE | SL | PyTorch | ✓ | |
IFED [162] | ECCV | Distortion | CNNs | Charbonnier, perceptual, TV loss | RS-GOPRO [162], FastecRS[55] | PSNR, SSIM, LPIPS | SL | PyTorch | ✓ | |
2023 | PM-Calib [163] | TPAMI | Standard | CNNs | Kullback-Leibler loss | 360Cities [164] | Mean error, Human sensitivity | SL | - | ✓ |
PerspectiveField [165] | CVPR | Standard | CNNs+Transformers | Cross-entropy loss | 360Cities [164], Stanford2D3D [166], TartanAir [150] | Mean error | SL | PyTorch | ✓ | |
Orienternet [167] | CVPR | Standard | CNNs | Log-likelihood loss | MGL [167] | Positions and rotation errors | SL | PyTorch | ||
Neumap [168] | CVPR | Standard | CNNs+Transformers | , cross-entropy loss | Cambridge Landmarks [71], ScanNet [88], 7scenes [148] | Median error | SL | PyTorch | ||
SESC [169] | IROS | Standard | CNNs | loss | KITTI [95], DDAD [170] | Median error | SSL | PyTorch | ||
CROSSFIRE [171] | ICCV | Standard | CNNs | , TV loss | Cambridge Landmarks [71], 7scenes [148] | Median error | SSL | PyTorch | ||
WildCamera [172] | NeurIPS | Standard | CNNs | Cosine similarity loss | ScanNet [88], MegaDepth [173], KITTI [95] | Median error | SL | PyTorch | ✓ | |
DroidCalib [174] | ICCV | Standard | CNNs | loss | TartanAir [150], EuRoC [141], TUM RGB-D [151] | Median error | SL | PyTorch | ✓ | |
NeuralLens [175] | CVPR | Distortion | CNNs | loss | SynLens [175] | RMS | SL | PyTorch | ✓ | |
DDA [176] | ICCV | Distortion | Diffusion | Diffusion loss | Places2 [98], Woodscape [177] | PSNR, SSIM, MS-SSIM, FID, LPIPS | USL | PyTorch | ✓ | |
CACM-Net [178] | CVPR | Distortion | CNNs | loss | Wireframes[101] | StraightAcc, ShapeAcc, ConformalAcc | SL | - | ✓ | |
RDTR [179] | TIP | Distortion | CNNs+Transformers | Cross-entropy, loss | Places2 [98] | EPE, PSNR, SSIM | SL | PyTorch | ✓ | |
DaFIR [180] | TCSVT | Distortion | Transformers | , loss | Places2 [98] | PSNR, SSIM | SL | PyTorch | ✓ | |
SimFIR [181] | ICCV | Distortion | Transformers | Cross-entropy, loss | Places2 [98] | PSNR, SSIM, FID | SL | PyTorch | ✓ | |
RecRecNet [182] | ICCV | Distortion | CNNs | , loss | MS-COCO[96] | PSNR, SSIM, FID, LPIPS | SL | PyTorch | ✓ | |
SDP-Net [183] | AAAI | Distortion | CNNs | , loss | DAVIS [184], Youtube-VOS [185] | PSNR, SSIM, FID, EPE | SL | PyTorch | ✓ | |
Darswin [186] | ICCV | Distortion | Transformers | Cross-entropy loss | ImageNet [76] | Accuracy | SL | PyTorch | ✓ | |
REG-Net [187] | ACM MM | Distortion | CNNs | loss | REG-HDR [187] | PSNR, SSIM, LPIPS | SL | PyTorch | ||
SelfDRSC [188] | ICCV | Distortion | CNNs | loss | RS-GOPRO [162] | PSNR, SSIM, LPIPS, NIQE, NRQM, PI | SSL | PyTorch | ✓ | |
SSL-RSC [189] | arXiv | Distortion | CNNs | loss | Gev-RS [47], Fastec-RS [55], ERS-VFI [189] | PSNR, SSIM, LPIPS | SSL | - | ✓ | |
EvShutter [190] | CVPR | Distortion | CNNs | loss | RS-ERGB [190], Fastec-RS [55] | PSNR, SSIM, LPIPS | SL | - | ||
SelfUnroll [191] | arXiv | Distortion | CNNs | , TV loss | RS-ERGB [190], Fastec-RS [55], Gev-RS [47], DRE [191] | PSNR, SSIM, LPIPS | SSL | PyTorch | ✓ | |
PatchNet [192] | WACV | Distortion | CNNs | , TV loss | BS-RSCD [137], Fastec-RS [55] | PSNR, SSIM, LPIPS | SL | PyTorch | ✓ | |
JAMNet [193] | CVPR | Distortion | CNNs+Transformers | , TV loss | Carla-RS [55], Fastec-RS [55], BS-RSCD [137] | PSNR, SSIM, LPIPS | SL | PyTorch | ✓ | |
QRSC [194] | ICCV | Distortion | Transformers | , loss | Carla-RS [55], Fastec-RS [55], BS-RSCD [137] | PSNR, SSIM, LPIPS | SL | PyTorch | ✓ | |
Deep_HM [195] | ICCV | Distortion | CNNs | loss | Carla-RS [55], RS-Homo [195] | PSNR, SSIM | SL | PyTorch | ✓ | |
2024 | DM-Calib [42] | arXiv | Standard | Diffusion | Diffusion loss | NuScenes [196], KITTI [95], SUN3D [197], CityScapes [198], etc | Relative error | SL | PyTorch | ✓ |
GeoCalib [51] | ECCV | Standard | CNNs | loss | OpenPano [51] | Median error, AUC | SL | PyTorch | ✓ | |
ExtremeRotation [199] | arXiv | Standard | Transformers | Cross-entropy loss | ExtremeLandmarkPairs [199] | Geodesic error | SL | PyTorch | ||
GAT-Calib [200] | WACV | Standard | GNNs | Cross-entropy, loss | World Cup 2014 [201] | IoU | SL | - | ✓ | |
SOFI [202] | BMVC | Standard | CNNs+Transformers | Focal loss | Google Street View [105], SUN360 [81], HLW[73] | Mean error, AUC | SL | PyTorch | ||
PWT-Calib [203] | WACV | Standard | CNNs | loss | CUTC [203] | Mean error, RMSE | SL | PyTorch | ||
NeFeS [204] | CVPR | Standard | CNNs+MLPs | , cosine similarity loss | Cambridge Landmarks [71], 7-Scenes [148] | Median error | SL | PyTorch | ||
U-ARE-ME [205] | arXiv | Standard | CNNs | loss | ICL-NUIM [206], TUM RGB-D [151], ScanNet [88] | ARE | SL | PyTorch | ✓ | |
FlowMap [207] | arXiv | Standard | CNNs | loss | MipNeRF-360 [208], Tanks & Temples [209], LLFF [210], CO3D [211] | ATE | SL | PyTorch | ||
MSCC [212] | WACV | Standard | CNNs+Transformers | Cross-entropy, angular distance loss | Google Street View [105], SYN-Citypark [54], Flickr [54], HLW [73] | Mean, median errors | SL | PyTorch | ✓ | |
HC-Net [213] | NeurIPS | Standard | CNNs | infoNCE loss | VIGOR [214], KITTI [95] | Mean, median errors | SL | Pytorch | ✓ | |
DiffCalib [41] | arXiv | Standard | Diffusion | Diffusion loss | Hypersim [215], NuScenes [196], KITTI [95], CitySpace [198], NYUv2 [108] | Relative error | SL | PyTorch | ✓ | |
CAR [40] | ICLR | Standard | Diffusion | Diffusion loss | CO3Dv2 [211] | Relative error | SL | PyTorch | ||
ADPs [216] | CVPR | Distortion | CNNs | loss | StreetLearn[132], SP360[161] | Mean absolute error, REPE | SL | PyTorch | ✓ | |
CDM [217] | TCSVT | Distortion | CNNs | , loss | MS-COCO [96], ADE20K [84], Wireframes [101] | PSNR, SSIM | SL | PyTorch | ✓ | |
QueryCDR [218] | ECCV | Distortion | CNNs+Transformers | loss | MS-COCO [96], Places2 [98] | PSNR, SSIM | SL | PyTorch | ✓ | |
VACR [219] | ECCV | Distortion | CNNs | loss | KITTI-360 [220], StreetLearn[132], Woodscape [177] | PSNR, SSIM, FID | SL | PyTorch | ✓ | |
DualPriorsCorrection [221] | ECCV | Distortion | CNNs+GANs | loss | Tan [53] | LineAcc, ShapeAcc | SL | PyTorch | ✓ | |
Disco [222] | IJCV | Distortion | GANs | loss | CMDP [223], USCPP [52] | LMK-E, PSNR, SSIM, LIPIPS, ID | SL | - | ||
MOWA [224] | arXiv | Distortion | Transformers | Cross-entroy, , loss | StitchRect [225], Place2 [98], MS-COCO [96], RotationCorr [226], Tan [53] | PSNR, SSIM, ShapeAcc | SL | PyTorch | ✓ | |
LBCNet [227] | TPAMI | Distortion | CNNs | loss | Carla-RS [55], Fastec-RS [55] | PSNR, SSIM | SL | PyTorch | ✓ | |
TACA-Net [228] | ACM MM | Distortion | Transformers | loss | Gev-RS [47], Fastec-RS [55] | PSNR, SSIM, LPIPS | SL | - | ✓ | |
UniINR [229] | ECCV | Distortion | CNNs | loss | Gev-RS [47], Fastec-RS [55] | PSNR, SSIM | SL | PyTorch | ✓ | |
DFRSC [230] | CVPR | Distortion | CNNs | loss | Carla-RS [55], Fastec-RS [55], BS-RSCD [137] | PSNR, SSIM, LPIPS | SL | PyTorch | ✓ | |
RS-Diffusion [43] | arXiv | Distortion | Diffusion | Diffusion loss | RS-Homo [195], RS-Real [43] | PSNR, SSIM, EPE | SL | PyTorch | ✓ |
2016 | DHN[29] | RSSW | Cross-View | VGG | loss | MS-COCO[96] | MSE | SL | Caffe | ✓ |
2018 | DeepFM[233] | ECCV | Cross-View | ResNet | loss | T&T[234], KITTI[95], 1DSfM[70] | F-score, Mean | SL | PyTorch | ✓ |
2019 | SSR-Net [65] | PRL | Cross-View | ResNet | loss | MS-COCO[96] | MAE | SSL | PyTorch | ✓ |
SRHEN [241] | ACM MM | Cross-View | CNNs | loss | MS-COCO [96], SUN397 [81] | MACE | SL | - | ✓ | |
@@
2021 | DLKFM [248] | CVPR | Cross-View | Siamese-Net | loss | MS-COCO[96], Google Earth, Google Map | MSE | SL | TensorFlow | ✓ |
@@
2022 | IHN [255] | CVPR | Cross-View | Siamese-Net | loss | MS-COCO[96], Google Earth, Google Map | MACE | SL | PyTorch | ✓ |
@@
Liu et al. [62] | TPAMI | Cross-View | ResNet | Triplet loss | Self-constructed | MSE, Accuracy | USL | PyTorch | ||
DXQ-Net [256] | IROS | Cross-Sensor | CNNs+RNNs | , geodesic loss | KITTI[95], KITTI-360[220] | MSE | SL | PyTorch | ✓ | |